OpenStudy (anonymous):

Express the required measurement in exact form (reduce fractions completely). What part of the circumference is an arc whose measure is a. 1°? c. 45°? e. 90°? b. 30°? d. 60°?

OpenStudy (anonymous):

We know there is 360 degrees in a circle So if we divide each by 360 and express as a reduced fraction we should be set a = 1/360, b = 45/360 = 1/8 and so on. OK?